### v2

# ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning

Abstract
Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility. This paper introduces ChatCFD, an automated agent system for OpenFOAM simulations that processes multi-modal inputs (e.g., research papers, meshes) via an interactive interface, leveraging DeepSeek-R1 and DeepSeek-V3 large language models (LLMs), a multi-agent architecture, and OpenFOAM knowledge. Its four-stage pipeline, Knowledge Base Construction, User Input Processing, Case File Generation, and Execution and Error Reflection, enables iterative trial-reflection-refinement for intricate setups, supporting diverse physical models and external meshes. Validation on 205 benchmark tutorial cases, 110 perturbed variants, and 2 literature-derived cases demonstrates ChatCFD’s 82.1% operational success rate on basic cases, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), and 60–80% on literature-derived complex cases, with turbulence model studies showing a 40% success rate for common models vs. 10% for rare ones like RNGk-$\epsilon$. Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors, such as dimensional inconsistency. Ablation studies highlight the efficacy of RAG-based modules and reflection mechanisms. By automating hypothesis testing and parameter exploration, ChatCFD accelerates scientific discovery in fluid mechanics and engineering, addressing LLM limitations through structured design and demonstrating strong potential as a modular component in MCP-based agent networks for collaborative multi-agent systems, paving the way for scalable AI-driven CFD innovation. The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
Keywords large language model, multi-agent system, computational fluid dynamics, OpenFOAM, automated CFD
1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [1, 2], energy systems [3, 4], urban environment [5, 6, 7], combustion [8, 9], and biomedical applications [10, 11]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [12, 13]. However, CFD implementation faces substantial barriers, requiring deep domain expertise [14] and often relying on expensive commercial software. Even experts invest significant time in tasks such as solver selection, model setup, mesh generation, boundary condition definition, and post-processing [15]. These demands, combined with high computational costs, limit CFD accessibility for smaller organizations and hinder broader innovation [13]. Thus, there is an urgent need for automated, intuitive, and affordable CFD solutions. To meet this need, we present ChatCFD, an AI-driven agent system that streamlines CFD workflows, facilitating automated scientific discovery in fluid mechanics and engineering by enabling rapid iteration and exploration of complex flow phenomena without extensive manual intervention.
Recent progress in Artificial Intelligence (AI) has transformed the automation of sophisticated scientific processes. Large Language Models (LLMs), such as GPT [16], Gemini [17], and DeepSeek [18], alongside multi-agent frameworks like MetaGPT [19] and AutoGen [20], excel in natural language processing, code generation, and reasoning via chain-of-thought techniques [21]. Frameworks like ReAct [22] synergize reasoning and acting in LLMs, while Reflexion [23] uses verbal reinforcement learning to improve agent performance through self-reflection. Toolformer [24] teaches LLMs to self-learn tool usage. Retrieval-Augmented Generation (RAG) [25] further bolsters these agents by incorporating domain-specific knowledge, reducing hallucinations, and improving domain adaptation for tasks like CFD automation.
This synergy has enabled LLM-based agents to automate CFD pipelines, from interpreting user inputs to running simulations. OpenFOAM [26], a leading open-source CFD platform, supports this trend due to its flexibility and community backing. Notable efforts include MetaOpenFOAM [27], which automates simulations from natural language using RAG and MetaGPT, though limited to tutorial-level cases and struggling with complex geometries; OpenFOAMGPT [28], evaluating RAG-augmented LLMs for zero-shot setup and boundary adjustments, but facing difficulties in turbulence model modifications; Foam-Agent [29], introducing hierarchical retrieval for dependency-aware generation, yet constrained in handling unseen configurations; and NL2FOAM [30], fine-tuning LLMs for natural language-to-CFD translation using LoRA [31], but reliant on domain-specific datasets and less effective for diverse physical models. Despite these advances, existing systems often fall short in end-to-end automation for complex cases without human oversight, particularly due to a fundamental challenge: LLMs lack sufficient training on niche scientific tasks like OpenFOAM setup, necessitating expert-designed architectures to bridge this gap through structured reasoning, reflection, and knowledge integration.
Existing CFD automation agents are often limited by their focus on rudimentary tasks, typically restricted to OpenFOAM tutorial-level examples, raising concerns about their generalization to complex, unseen cases. This challenge primarily stems from the scarcity of domain-specific scientific coding corpora, such as OpenFOAM setups, requiring sophisticated multi-agent system designs to effectively incorporate physical knowledge into agent architectures, thereby enhancing LLMs’ capabilities in specialized scientific tool invocation. Recent representative work in this area includes SciToolAgent [32], which develops a knowledge graph-driven agent for integrating hundreds of scientific tools across biology, chemistry, and materials science, using graph-based RAG and safety-checking to automate workflows like protein engineering and chemical synthesis. While not directly focused on CFD, its approach to tool integration and domain adaptation offers valuable insights for fluid automation agents in handling coupled physics and complex simulations.
Three critical challenges in CFD automation agents remain unaddressed. First, incorporating domain-specific knowledge into CFD agents remains challenging. Engineering-scale simulations require comprehensive initial and boundary condition specifications, necessitating domain-specific structural thinking and seamless integration of specialized domain expertise into the automation pipeline. Second, designing effective agent frameworks, especially for future multi-MCP agent collaborations, is essential to maximize complementary capabilities. Typical issues include whether we need to contain mesh generation in the agent scope, where real-world geometries exceed basic OpenFOAM tools like blockMesh and require specialized integration of external meshes. Third, enabling large-scale testing and evolution of fluid agents is crucial. Current reliance on simplistic natural language descriptions often leads to ambiguous case setups and hinders scalability. In contrast, CFD literature, comprising millions of papers with detailed descriptions and standard results, remains underutilized, impeding agent evolution. These shortcomings highlight a critical gap in current CFD automation frameworks.
Figure 1: Overview of the ChatCFD automated agent system for streamlining CFD simulations within the OpenFOAM framework. ChatCFD enables researchers and engineers to configure and execute simulations with minimal CFD or OpenFOAM expertise. The system comprises three core components: (1) an interactive chat interface for users to input case descriptions or upload mesh files, (2) a thinking system, the core decision-making module (detailed in Figure 2), and (3) a simulation engine that executes cases, collects error logs, and delivers final results.
To overcome these challenges, we introduce ChatCFD, an automated agent system for OpenFOAM simulations, as shown in Figure 1. ChatCFD processes multi-modal inputs (e.g., research papers, meshes) through an interactive interface, utilizing LLMs (DeepSeek-R1, DeepSeek-V3), multi-agent architecture, and OpenFOAM knowledge to manage the full workflow. It supports intricate setups via iterative refinement, handling diverse physical models and external meshes, surpassing prior agents in adaptability to unseen cases.
This paper details ChatCFD’s architecture and validates its components experimentally. The structure is as follows: Section 2 describes the pipeline design, Section 3 presents results, and Section 4 summarizes findings. Appendices provide examples and prompts for operational clarity.
2 ChatCFD Pipeline Design
Figure 2: Architecture of the ChatCFD framework for automated CFD simulations, illustrating the four-stage workflow and agent structure. The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Generation, generating OpenFOAM case files using the knowledge base; and (4) Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever. The agent structure integrates DeepSeek-R1 and DeepSeek-V3 for intelligent processing, with iterative error correction.
ChatCFD is an automated CFD agent system that leverages the OpenFOAM framework to process multi-modal user inputs, including research articles and mesh files, to configure and execute CFD simulations based on user instructions. As illustrated in Fig. 2, the ChatCFD framework implements a comprehensive workflow to streamline simulation setup, execution, and analysis.
2.1 Stage 0: Knowledge Base Construction
The preprocessing aims to establish a foundational knowledge base for CFD tasks, performed in advance to optimize ChatCFD’s operations. Ideally, this knowledge base contains general CFD principles (e.g., numerical differentiation, discretization schemes), OpenFOAM-specific manuals, and a comprehensive example library. Fine-tuning large language models (LLMs) for general CFD knowledge is beyond the current scope; readers are referred to recent work by Dong et al. [30]. This study focuses on preprocessing OpenFOAM manuals and publicly accessible examples. A systematic analysis of OpenFOAM manuals and tutorials constructs a structured JSON database, cataloging essential parameters such as solver configurations, turbulence models, and file dependencies, providing a critical foundation for accurate case setup and error correction.
Initially, preprocessed manual and tutorial data are used to define file dependency relationships, which are accessed during the RAG process to enhance the agent’s error correction capabilities. To ensure robust operation, strategic case filtering is implemented by removing OpenFOAM-specific headers and excluding cases with external dependencies or auxiliary folders. This maintains case independence, reducing the risk of degraded LLM response quality due to irrelevant configuration elements. For cases where non-uniform field definitions occupy significant file content and may exceed the dialogue context limits of large models, placeholder tokens such as __CELL_COUNT__,__FACE_COUNT__,__SCALAR_DATA_PLACEHOLDER__ are used to replace non-uniform field data. This approach enables the LLM to focus on critical settings, such as boundary and initial conditions, while ignoring non-uniform field configurations. As a result, it mitigates the reduction in the number of usable OpenFOAM tutorial cases caused by complex non-uniform field data. Filtering criteria and implementation details are provided in Appendix A. Subsequently, the example library is preprocessed through systematic extraction and organization. All case configuration files are parsed and converted into a structured JSON format, tagged with metadata such as solver specifications, turbulence model types, and flow regime characteristics. Analysis of solver and turbulence model distributions determines specific file requirements for each configuration. For instance, the simpleFoam solver requires configuration files including system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, constant/transportProperties, and constant/turbulenceProperties. Similarly, the k-$\omega$ SST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
2.2 Stage 1: User Input Processing
The initial stage of ChatCFD features an interactive, multi-modal interface that enables users to define CFD simulations by either conversing with the DeepSeek-R1 model or uploading documents and mesh files, built using the Streamlit Python framework. This interface leverages the preprocessed knowledge base to facilitate accurate case specification with minimal CFD expertise. The knowledge base, comprising OpenFOAM manuals, tutorial cases, and a structured JSON database of solver configurations and turbulence models, informs the system’s natural language processing and case extraction, ensuring robust and context-aware interactions. The workflow comprises four key phases:
  • Input Submission: Users can upload research articles or describe the case through dialogue to provide the basis for case analysis.
  • Case Extraction: The system extracts and catalogs CFD cases from uploaded documents or conversation inputs, identifying solver configurations and physical models, and presents them in a structured format with unique identifiers (e.g., Case 1, Case 2).
  • Case Selection: Through natural language interaction, the system guides users to select a target case, provides detailed specifications for verification, and confirms the selection.
  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent format .msh files. For OpenFOAM default meshes, users must specify the path to the corresponding polyMesh directory.
Upon completing this workflow, ChatCFD aggregates three essential components for downstream processing, as shown in Fig. 2:
  • Case Documentation: The system retains the full research article or conversation details, providing access to all technical specifications, including solvers, boundary conditions, and simulation parameters.
  • Case Specification: The selected case is recorded with detailed metadata, including solver types, numerical schemes, physical models, and source references, ensuring precise setup.
  • Mesh Data: Robust file transfer and path management ensure reliable delivery of mesh data to subsequent stages.
2.3 Stage 2: Case File Generation
The CFD engineer layer, powered by the DeepSeek-R1 and DeepSeek-V3 models, initializes OpenFOAM case files through a streamlined three-phase process that leverages the preprocessed knowledge base.
2.3.1 Phase 1: File Configuration Synthesis
The system analyzes the case description from Stage 1 (User Input Processing) to identify suitable solvers and models, thereby generating a list of required case file configurations. This involves explicitly matching dependencies for solvers, models, and related settings to facilitate accurate file configuration. The workflow proceeds as follows:
  • Relevant cases are retrieved based on the solver and model, yielding a corresponding file configuration list from which non-essential files (e.g., those for visualization or mesh generation) are removed. These refined cases are then input to the DeepSeek-V3 model to produce an optimized file configuration list.
  • In the absence of direct matches, cases are aligned by solver type. For additional models, such as turbulence models, necessary files are inferred from analogous cases, synthesized into a new configuration list, verified for solver compatibility, and finalized via DeepSeek-V3.
This approach integrates solver and physical model structures with case synthesis, enhancing initialization precision, minimizing file omissions, and substantially improving overall case file quality.initialization.
2.3.2 Phase 2: Case Setup Extraction
Detailed case configurations are derived from three primary sources: the file configuration list, user prompts, and the case library. The configuration list prevents omissions or redundancies in the case structure, while user prompts enable filtering of segmented case description chunks using the sentence-transformer/all-mpnet-base-v2 library and a similarity threshold. The DeepSeek-R1 model implements a hierarchical three-step extraction:
  • Mesh Boundary Condition Extraction: Boundary names and types are extracted via regular expressions and tools like pyFoam, forming a persistent dictionary for subsequent physical field generation and error handling.
  • Boundary Condition Validation: Boundary conditions are identified from filtered chunks, checked for compatibility with OpenFOAM-v2406, and subjected to rigorous naming and format validation.
  • Physical Field File Value Setup: Field files (e.g., 0/p, 0/nut, 0/nutTilda) are identified from the configuration list. Using prior boundary data, OpenFOAM-compliant templates are created, encompassing internal and boundary field settings differentiated by boundary names and associated keywords.
2.3.3 Phase 3: Configuration Validation and Refinement
For unspecified parameters, such as discretization schemes and solver algorithms in system/fvSolution and system/fvSchemes, the DeepSeek-R1 model applies advanced reasoning grounded in CFD best practices and case-specific physics to formulate coherent configurations. This maintains parameter interdependencies and ensures alignment with the outlined flow dynamics and solver demands.
A comprehensive validation and correction mechanism further refines the generated files. The system scrutinizes structures in the system and constant directories, confirming proper dependencies and completeness through cross-references with library exemplars. Particular emphasis is placed on physical field dimensions to mitigate LLM limitations in interpreting such constraints. For example, variables like p and alphat—which differ dimensionally between compressible and incompressible flows—are calibrated according to case type; ambiguous cases retain initial values. This process curtails errors like dimensional inconsistencies, bolsters initial file robustness, and diminishes iterative corrections in Stage 3.
2.4 Stage 3: Error Correction and Reflection
Stage 3 of the ChatCFD pipeline employs a modular architecture for automated error correction and iterative refinement in OpenFOAM case configurations. The following modules facilitate robust error handling:
  • ReferenceRetriever: This module retrieves reference files from the preprocessed OpenFOAM knowledge base, matching solver and model specifications from benchmark tutorial cases. If no exact match is found, it prioritizes solver-compatible files. To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error localization and resolution.
  • ContextRetriever: This module compiles current case configurations into a structured JSON format, providing detailed file content and modification trajectories over a specified period, tailored to other modules’ needs. It supports targeted reflection and correction, particularly for errors involving physical coupling (e.g., pressure-density interactions).
  • Error Localization Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise localization.
  • Error Correction Module: This module addresses file interdependencies in OpenFOAM through a three-step process: (1) DeepSeek-V3 identifies files requiring coordinated modifications to resolve coupling-related errors; (2) DeepSeek-R1 proposes actionable corrections by integrating error details, related file content, simulation requirements, and benchmark tutorial cases; (3) DeepSeek-V3 applies these corrections, ensuring compliance with OpenFOAM’s file format standards.
  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with localization and correction modules.
Errors detected during simulation are categorized into two types:
  • General Errors: These, comprising over 70% of runtime errors, include incorrect keywords, formatting issues, or floating-point errors. Correction involves a streamlined workflow: suspicious file detection (DeepSeek-V3), error confirmation (DeepSeek-R1), retrieval of related tutorial files (ReferenceRetriever), proposal of corrections, and file modification. For missing file errors, DeepSeek-V3 directly generates the required file using reference cases, followed by dimensional validation to ensure quality.
  • Persistent Errors: ChatCFD leverages short-term memory (recent error messages and file modification histories) and long-term memory (reflection histories) to address recurring issues. Reflection histories, stored as structured insights, enhance the system’s ability to adapt and resolve complex errors iteratively.
2.5 LLM Architecture and Configuration
ChatCFD employs a dual-model architecture, integrating the DeepSeek-R1 and DeepSeek-V3 large language models (LLMs) to execute a comprehensive CFD workflow for case generation and error correction. Each model is strategically deployed based on its strengths to optimize performance across Stages 2 and 3.
The DeepSeek-R1 model, with its superior text comprehension and reasoning capabilities, is utilized for complex, high-stakes tasks. In Stage 2, it extracts critical simulation parameters (e.g., boundary conditions, initial conditions) from case descriptions and generates complete initial case files. In Stage 3, it pinpoints error-causing files, proposes detailed correction strategies, and conducts reflection on persistent errors. However, DeepSeek-R1’s propensity for hallucinations—such as inserting unnecessary Markdown formatting, which can disrupt error correction—poses challenges. Despite mitigation efforts using prompt engineering and pydantic for structured outputs, these issues persist. Additionally, its inference time increases significantly with longer prompts, making it less efficient for routine tasks.
The DeepSeek-V3 model excels in instruction following and rapid response, making it ideal for structured, low-complexity tasks. In Stage 2, it performs quick file structure validation, while in Stage 3, it identifies suspicious or missing files and generates corrected file content based on provided instructions. To address DeepSeek-V3’s limited reasoning capacity, tasks are simplified into structured formats (e.g., providing configuration file lists to narrow analysis scope), and the model is prompted to justify its actions (e.g., explaining why a file is suspicious). This enhances its effectiveness while maintaining computational efficiency. This dual-model approach leverages DeepSeek-R1’s reasoning for complex CFD tasks and DeepSeek-V3’s efficiency for structured operations, ensuring robust performance across diverse case types (e.g., benchmark tutorial cases, literature-derived cases) while minimizing computational costs.
ChatCFD maintains a comprehensive log for each CFD case, capturing all question-answer interactions and actions taken during reflection-based error correction. These logs are invaluable for analyzing the behavior of the DeepSeek-R1 and DeepSeek-V3 models within the OpenFOAM framework, particularly in identifying limitations in handling complex CFD configurations. The recorded reflection iterations, which document corrective actions for persistent errors, provide critical insights into scenarios where ChatCFD’s performance is suboptimal, guiding targeted improvements in the pipeline. Additionally, a concise summary of reflections for recurring errors is maintained, offering a deeper understanding of ChatCFD’s behavior in complex scenarios. These logs have the potential to form a novel database, serving as a valuable resource for enhancing LLM-based agents in addressing high-complexity CFD problems within the MCP framework.
The experimental validations described in Section 3 were conducted on the VolcEngine platform [33], which provides the pricing structure for token usage and computational costs. For DeepSeek-R1, input tokens cost \$0.00055 per 1,000 (0.004 RMB), and output tokens cost \$0.0022 per 1,000 (0.016 RMB). For DeepSeek-V3, input tokens cost \$0.00021 per 1,000 (0.0015 RMB), and output tokens cost \$0.00082 per 1,000 (0.006 RMB). All token consumption and cost metrics reported in Section 3 are based on this pricing, enabling precise evaluation of ChatCFD’s computational efficiency across benchmark tutorial cases, perturbed variant cases, and literature-derived cases.
3 Results and discussion
The performance of ChatCFD was rigorously evaluated through a series of validation experiments encompassing three distinct categories of CFD cases: (i) 205 benchmark tutorial cases drawn from OpenFOAM tutorials and the OpenFOAM wiki, serving as standardized references for foundational validation; (ii) 110 perturbed variants, derived by systematically altering key parameters (e.g., boundary conditions, solver settings, or physical properties) in the benchmark cases to assess robustness and sensitivity; and (iii) 2 advanced literature-derived cases, directly prompted from published research articles to test real-world applicability in complex, unseen scenarios. ChatCFD demonstrated an operational success rate of 82.1% across the 315 benchmark and perturbed cases, with success defined as error-free configuration and execution leading to converged simulations. For the advanced literature-derived cases, success rates ranged from 60% to 80%, reflecting the challenges of handling intricate, domain-specific configurations without prior exposure. These results represent a substantial advancement in LLM-driven CFD automation, particularly in reducing the expertise barrier for setup and execution within the OpenFOAM framework.
3.1 Performance Evaluation on OpenFOAM Basic Cases
Figure 3: Comparison of success rates across three CFD agents (ChatCFD, MetaOpenFOAM, Foam-Agent) for 205 benchmark and 110 perturbed OpenFOAM tutorial cases. (a) Success rates by case category. (b) Distribution of test cases across categories. (c) Overall success rate comparison.
To evaluate ChatCFD, a comprehensive set of test cases was curated from OpenFOAM tutorials and the OpenFOAM wiki. Cases involving complex mesh generation techniques were filtered out, resulting in 205 benchmark tutorial cases. From these, 11 cases were selected and systematically perturbed by modifying their descriptions and adjusting parameters (e.g., boundary conditions, solver settings, or physical properties), yielding an additional 110 perturbed variant cases. This approach ensures both foundational validation and robustness testing under varied conditions. Figure 3 illustrates the performance of ChatCFD compared to MetaOpenFOAM and Foam-Agent across these cases, showing success rates by category, case distribution, and overall performance. Success is defined as the generation of error-free case files leading to converged simulations. As shown in Figure 3(c), ChatCFD achieves an overall success rate of 82.1%, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%). Figure 3(a) highlights ChatCFD’s consistent advantage across case categories, demonstrating its superior performance and generalization capability.
[…]
Figure 4: Performance statistics for different agents across 205 benchmark tutorial cases. (a) Average token consumption per case. (b) Distribution of token consumption. (c) Average number of reflection iterations per case. (d) Average token consumption per reflection iteration. (e) Average computational cost (in monetary terms). (f) Distribution of reflection iteration ratios, excluding zero and limit-reaching cases.
The performance of ChatCFD, MetaOpenFOAM, and Foam-Agent was evaluated through detailed metrics on 205 benchmark tutorial cases from OpenFOAM, focusing on token consumption, reflection iterations, and computational cost. Figure 4 presents a comparative analysis of these metrics, highlighting ChatCFD’s efficiency and robustness in automated CFD workflows. As shown in Figure 4(a), ChatCFD exhibits the lowest average token consumption per case, approximately half that of Foam-Agent, with MetaOpenFOAM ranking second. This efficiency underscores ChatCFD’s high success rate combined with reduced costs, making it practical for large-scale CFD applications. While ChatCFD and MetaOpenFOAM have comparable total token consumption, Figure 4(e) reveals a significant cost disparity, with ChatCFD’s computational cost being approximately 1.5 times lower than MetaOpenFOAM’s. This is primarily due to ChatCFD’s strategic use of the cost-efficient DeepSeek-V3 model (with a cost half that of DeepSeek-R1) for specific pipeline stages, as illustrated in Figure 4(b). In contrast, MetaOpenFOAM’s lower success rate necessitates frequent error corrections, generating additional file content and inflating costs. Figure 4(c) shows that MetaOpenFOAM requires the highest average number of reflection iterations per case, followed by Foam-Agent, with ChatCFD requiring the fewest. However, Figure 4(d) indicates that MetaOpenFOAM consumes fewer tokens per reflection iteration due to its simpler error-correction approach, which dilutes the high token cost of initial case setup across multiple iterations. Conversely, ChatCFD and Foam-Agent employ more sophisticated error-reflection strategies. Foam-Agent, for instance, maintains an “error correction trajectory” that logs modifications to file $i$ at iteration $j$, supplemented by contextual data from related files and reference cases. This increases its per-iteration token consumption. ChatCFD, while similarly providing contextual data, optimizes efficiency by maintaining both short-term modification trajectories and long-term reflection summaries. This dual approach enables precise filtering of irrelevant information and concise summarization of critical context, significantly reducing token consumption during LLM interactions compared to Foam-Agent, as evidenced in Figure 4(d). Figure 4(f) depicts the distribution of reflection iteration ratios, defined as the number of reflections divided by the maximum allowed iterations, reflecting each agent’s assessment of case complexity. The x-axis represents the reflection ratio, and the y-axis indicates the proportion of cases at each complexity level. Higher curves indicate stronger error-correction capabilities at a given reflection ratio. All three agents exhibit a rapid decline in solved cases followed by saturation, suggesting that increasing reflection iterations beyond a certain point yields diminishing returns. For Foam-Agent, with a reflection limit of 49 iterations, the curve flattens at approximately 50% (around 25 iterations), indicating that additional reflections beyond this threshold rarely resolve errors. ChatCFD’s curve consistently lies above the others, demonstrating more effective reflections that are likelier to correct errors. However, even for ChatCFD, reflections beyond 75% of the limit (approximately 22 iterations) often fail to resolve remaining errors, highlighting a current limitation in LLM-based CFD agents where complex case errors require advanced knowledge integration or alternative strategies beyond iterative reflection.
3.2 Evaluation and Analysis on Literature-Derived CFD Cases
Figure 5: Visualization of two literature-derived CFD cases. (a) Computational mesh for the NACA0012 airfoil case [34]. (b) Computational mesh for the Nozzle case [35]. (c) Boundary condition types for the NACA0012 case. (d) Boundary condition types for the Nozzle case.
To evaluate ChatCFD’s ability to handle complex, real-world scenarios, two representative literature-derived cases were selected: an incompressible flow case and a compressible flow case, as illustrated in Figure 5. These cases were chosen for their comprehensive documentation in the source literature, providing detailed specifications for solver configurations, turbulence models, and mesh characteristics critical for accurate CFD reproduction. The incompressible flow case, based on Sun et al. [34], involves a NACA0012 airfoil at a 10° angle of attack, simulated using the simpleFoam solver with the Spalart-Allmaras turbulence model. The compressible flow case, drawn from Yu et al. [35], models a nozzle with a pressure ratio of 3, employing the rhoCentralFoam solver and the Spalart-Allmaras model.
Given reported challenges in modifying OpenFOAM turbulence models using large language models (LLMs), as noted by Pandey et al. [28], additional experiments were conducted to explore robustness across various turbulence models (e.g., $\epsilon, k$-$\omega$ SST). Table 1 provides a comprehensive summary of all case configurations and experimental iterations. High-fidelity computational meshes were generated to ensure accuracy. The experimental protocol was structured as follows: Cases 1 (NACA0012) and 5 (nozzle) underwent an extensive model ablation study, with 10 iterations per case across five system configurations to assess sensitivity to model variations. Cases 2 through 4, using different turbulence models, were evaluated based on the full ChatCFD system configuration, with 10 iterations each, to focus on performance under optimal settings. This rigorous methodology resulted in 130 experimental runs, forming a robust dataset for analyzing ChatCFD’s performance in complex, literature-derived scenarios. The results, detailed in subsequent analyses, demonstrate ChatCFD’s capability to adapt to sophisticated CFD configurations, complementing its strong performance on benchmark tutorial and perturbed variant cases.
Table 1: Summary of Test Cases for ChatCFD Validation
Case ID Case Mesh Physical Models Experimental Design
Case 1 NACA0012
simpleFoam with
Spalart-Allmaras
10 runs $\times$ 5 system
configurations
Case 2 NACA0012
simpleFoam with
k-$\omega$ SST
10 runs $\times$ the complete
system configuration
Case 3 NACA0012
simpleFoam with
k-$\epsilon$
10 runs $\times$ the complete
system configuration
Case 4 NACA0012
simpleFoam with
RNGk-$\epsilon$
10 runs $\times$ the complete
system configuration
Case 5 Nozzle
rhoCentralFoam with
hePsiThermo and
Spalart-Allmaras
10 runs $\times$ 5 system
configurations
3.2.1 Ablation Study on ChatCFD System Configurations
To systematically dissect the contributions of ChatCFD’s core components, ablation experiments were conducted across five distinct system configurations. These configurations vary in the complexity of case file generation (Stage 2, Section 2.3), the error-handling logic, and the integration of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files). The configurations are defined as follows:
  • Configuration A (Baseline System): Employs a simplified parameter extraction process in Stage 2, bypassing the three-step hierarchical extraction. In Stage 3, error correction is limited to a basic ’general error’ pathway without ReferenceRetriever (tutorial-based retrieval).
  • Configuration B (Baseline with ReferenceRetriever) and Simplified Setup Extraction: Retains the simplified Stage 2 extraction but enhances Stage 3 with the ReferenceRetriever module, enabling tutorial-based error correction within the general error pathway.
  • Configuration C (Enhanced Setup and Reference Retrieval): Implements the full three-step setup extraction and configuration validation in Stage 2. Stage 3 includes general error correction augmented by the ReferenceRetriever module for improved accuracy.
  • Configuration D (Full Error Reflection with Simplified Extraction): Uses the simplified Stage 2 extraction but incorporates both general and persistent error correction modules in Stage 3, supported by ReferenceRetriever and ContextRetriever for comprehensive error handling.
  • Configuration E (Complete System): Represents the full ChatCFD pipeline, featuring comprehensive case setup extraction and validation in Stage 2 and complete error reflection modules in Stage 3, integrating both ReferenceRetriever and ContextRetriever for optimal performance.
Figure 6: Success rates of ChatCFD across five system configurations (A–E) for literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5). The horizontal axis represents configurations, and the vertical axis quantifies success rates. The 10-step operational success rate indicates completion of 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. The accurate configuration success rate reflects precise adherence to source article specifications.
Figure 6 illustrates ChatCFD’s performance across Configurations A through E for two literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5). Two metrics were evaluated: (1) the 10-step operational success rate, defined as the ability to complete 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and (2) the accurate configuration success rate, indicating exact compliance with the solver, turbulence model, and boundary condition specifications in the source articles. The 10-step criterion enables rapid assessment of operational stability, particularly for the reflection mechanism, as literature often lacks specified simulation durations. Configurations achieving this milestone were generally free of major operational faults, though minor deviations in boundary condition parameters could persist. Cases meeting the accurate configuration criterion exhibited simulation results in strong agreement with published data, as detailed in the appendix. These results underscore ChatCFD’s robustness in handling complex, literature-derived configurations, complementing its performance on benchmark tutorial and perturbed variant cases.
For the NACA0012 Case 1, the data in Figure 6 show a clear progression in the 10-step operational success rate with increasing system sophistication. The baseline Configuration A achieved a 20% success rate. This trend highlights the positive impact of incrementally integrating retrieval modules and OpenFOAM-specific knowledge. However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, unable to precisely replicate the case setup from the source article. This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and setup extraction as part of its advanced Stage 2 processing. With Configuration E, the accurate configuration success rate for the NACA0012 case improved significantly to 40%, while maintaining the 80% operational success rate. Analysis of the DeepSeek-R1 model’s internal reasoning revealed that, without robust article interpretation and strict compliance with documented specifications in Stage 2, the LLM often defaults to generic boundary conditions and simplified parameters, such as standard inlet-type conditions for inflow velocities or pressureOutlet-type conditions for outflow pressures, instead of the specialized freestream conditions required for the NACA0012 airfoil. These erroneous configurations underscore the critical role of comprehensive case setup extraction and validation in Stage 2, which Configuration E leverages to ensure fidelity to complex CFD requirements, enhancing ChatCFD’s performance in complicated scenarios.
For the Nozzle Case 5, Figure 6 demonstrates a significant enhancement in the 10-step operational success rate with the integration of the ContextRetriever module. Configurations A through C, lacking ContextRetriever, achieved a 0% success rate for this metric, defined as completing 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. Configuration D, incorporating ContextRetriever, improved this rate to 60%. This advancement stems from ContextRetriever’s ability to diagnose complex coupling-related errors within the MCP framework, addressing a key limitation of earlier configurations. For example, errors in thermophysical parameters, such as compressibility ($\Psi$), were often misattributed to the constant/thermophysicalProperties file in prior configurations. In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T). By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error localization and effective correction, enhancing performance for complex cases.
The differential impact of ContextRetriever between the NACA0012 Case 1 and Nozzle Case 5 reflects their distinct flow physics. The NACA0012 case, an incompressible flow simulation using the pisoFoam solver, exhibits weak pressure-velocity coupling, reducing the need for advanced error handling. Conversely, Nozzle Case 5, a compressible flow scenario with the rhoCentralFoam solver, involves strong interdependencies among density, temperature, and pressure, governed by the equation of state. The ContextRetriever module’s advanced analysis of inter-file dependencies proves particularly effective for the Nozzle case’s coupled physics, yielding greater performance gains compared to the less coupled NACA0012 case. This highlights ContextRetriever’s proficiency in managing complex physical interactions across diverse CFD scenarios.
Advancing to Configuration E, which integrates comprehensive article interpretation in Stage 2, maintained the 60% 10-step operational success rate for Nozzle Case 5 while increasing the accurate configuration success rate from 0% to 30%. This metric reflects precise adherence to source specifications. Consistent with findings for the NACA0012 case, Configuration E’s enhanced setup extraction and validation ensure greater configuration fidelity across flow regimes, reinforcing ChatCFD’s robustness in literature-derived cases.
Figure 7: ChatCFD’s performance metrics for accurate configuration of literature-derived NACA0012 Case 1 and Nozzle Case 5. Metrics include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection).
The performance disparity between the literature-derived NACA0012 Case 1 and Nozzle Case 5 arises from two key factors: (1) the increased complexity of managing numerous files with intricate content in compressible flows, and (2) the stringent requirement for consistency across coupled thermophysical models (e.g., 0/p, 0/T, constant/thermophysicalProperties). The Nozzle case, a compressible flow simulation using rhoCentralFoam, involves strong coupling among pressure, density, temperature, and the equation of state, necessitating precise dimensional and physical alignment across governing equations and material properties. This contrasts with the NACA0012 case, an incompressible flow simulation using pisoFoam with weaker pressure-velocity coupling, which imposes fewer constraints. Additionally, an implicit bias in LLM training data, predominantly focused on incompressible flow scenarios, leads to erroneous defaults in compressible cases. For instance, in OpenFOAM, incompressible flows use kinematic pressure ($p_{k}$, dimensions $L^{2}T^{-2}$)—dynamic pressure divided by density—while compressible flows require absolute thermodynamic pressure ($p$, dimensions $ML^{-1}T^{-2}$). The DeepSeek-R1 model’s tendency to apply incompressible conventions to compressible cases introduces dimensional errors, such as incorrect pressure settings, requiring additional reflection iterations for correction.
Figure 7 quantifies these challenges: the Nozzle case requires over twice the reflection iterations, triple the DeepSeek-R1 tokens, and six times the DeepSeek-V3 tokens compared to the NACA0012 case, resulting in approximately triple the execution cost. The disproportionate DeepSeek-V3 token increase (6x vs. 3x for DeepSeek-R1) reflects frequent dimensional inconsistencies and persistent errors, necessitating more invocations of the costlier DeepSeek-V3 model. Figure 7(c) confirms this, showing a DeepSeek-V3 token multiplier of  2.7x per reflection for the Nozzle case, compared to  2x for DeepSeek-R1, relative to the NACA0012 case. These metrics highlight the computational burden of complex flow simulations and underscore ChatCFD’s ability through advanced error correction.
3.2.2 Impact of Turbulence Models on ChatCFD Performance
Figure 8 presents ChatCFD’s performance across four turbulence models—Spalart-Allmaras (SA), $\epsilon, k$-$\omega$ SST, and RNG k-$\epsilon$—applied to the literature-derived NACA0012 cases (Cases 1 to 4). The analysis focuses on simulations achieving accurate configuration per published specifications, excluding those with only basic operational success (e.g., 10-step operational success rate). All results were obtained using the complete ChatCFD system (Configuration E).
As shown in Figure 8(a), success rates for accurate configuration vary significantly: SA at 40%, k-$\omega$ SST at 30%, k-$\epsilon$ at 20%, and RNG k-$\epsilon$ at 10%. This demonstrates ChatCFD’s ability to handle diverse turbulence models, addressing LLM challenges noted in OpenFOAMGPT [28] through structured knowledge integration and RAG. Performance metrics in Figure 8(a,b) reveal consistent profiles for SA, k-$\epsilon$, and k-$\omega$ SST, with an average of  6 reflection iterations, comparable DeepSeek-R1 token usage, and costs. In contrast, RNG k-$\epsilon$ requires  23.5 reflections—nearly 4x higher—elevating token consumption and costs.
Figure 8(c) attributes these disparities to configuration file token counts and model prevalence in OpenFOAM tutorials. While file sizes differ modestly ( 1,800 tokens for SA vs.  2,000 for others), tutorial distribution is more influential: SA, k-$\epsilon$, and k-$\omega$ SST appear in 28+ tutorials each, versus one for RNG k-$\epsilon$. RNG k-$\epsilon$’s performance deficit stems from: (1) its rarity in engineering applications, leading to sparse LLM training data and poor zero-shot configuration; and (2) limited ReferenceRetriever efficacy due to few OpenFOAM tutorials, hindering guidance for parameters in constant/turbulenceProperties, system/fvSchemes, and system/fvOption. This necessitates greater reliance on the LLM’s limited intrinsic knowledge, reducing performance relative to common models.
Figure 8: ChatCFD’s average performance for accurate configuration of NACA0012 Cases 1 to 4 across turbulence models (Spalart-Allmaras, $\epsilon, k$-$\omega$ SST, RNG k-$\epsilon$). Metrics shown include: accuracy (success rate of precise configuration according to literature), number of reflection iterations (#reflection), total token consumption of the DeepSeek-R1 model (R1 total token), execution cost (Cost \$), token count of configuration files (#token of config files), and the number of cases incorporating each turbulence model in OpenFOAM tutorials (#Case of model in OF tutorial).
3.3 Future Potential of ChatCFD in Multi-Agent CFD Workflows
Building on the insights from ChatCFD’s performance in benchmark tutorial, perturbed variant, and literature-derived cases, ongoing work explores its potential as a foundational component in advanced engineering workflows. Beyond functioning as a standalone tool for CFD automation, ChatCFD is being adapted through the Model Context Protocol (MCP), a standardized framework for integrating specialized AI agents into a cohesive multi-agent system. MCP facilitates seamless communication and data exchange among agents by defining structured protocols for context sharing, enabling ChatCFD to collaborate with other intelligent tools to address complex, interdisciplinary engineering tasks.
As illustrated in Figure 9, this multi-agent workflow enables an end-to-end design cycle for CFD applications. The process begins with a 3D modeling agent that generates a three-dimensional geometry from a natural language description or a simple image, such as a sketch or photograph. A subsequent meshing agent automatically discretizes the geometry into a high-fidelity computational mesh, leveraging tools like commercial mesh generators or open-source alternatives (e.g., blockMesh, snappyHexMesh). This mesh is then passed to the ChatCFD agent, which autonomously configures solver parameters, boundary conditions, and turbulence models (e.g., simpleFoam with Spalart-Allmaras or rhoCentralFoam with k-$\omega$ SST) to execute the CFD simulation, computing key performance metrics such as drag coefficients or pressure distributions. Finally, an optimization agent analyzes the simulation results, providing insights for iterative design refinement, such as adjusting geometry or boundary conditions to minimize drag or enhance flow stability. This MCP agent is available to use and can be found online (https://www.bohrium.com/apps/designagent0812).
This integrated approach highlights ChatCFD’s transformative potential beyond simulation automation. By serving as a standardized computational module within a collaborative multi-agent ecosystem, ChatCFD can contribute to broader engineering design processes, from conceptual design to performance optimization. The MCP framework ensures interoperability, allowing ChatCFD to adapt to diverse tasks, such as aerodynamics, heat transfer, or multiphase flow simulations, by interfacing with agents specialized in geometry generation, mesh refinement, or post-processing. This vision positions ChatCFD as a cornerstone for future AI-driven engineering workflows, paving the way for scalable, interdisciplinary applications in scientific discovery and industrial innovation.
Figure 9: ChatCFD integrated within a multi-agent workflow via the Model Context Protocol (MCP), enabling transformation from a single sentence or image to a complete CFD simulation.
4 Conclusion
This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations. By incorporating a multi-agent architecture, domain-specific knowledge bases, and advanced error reflection mechanisms, ChatCFD achieves end-to-end automation for reproducing literature results, including both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flows. The simulations accurately reproduce velocity contours, pressure coefficients, and schlieren images, aligning closely with experimental data. ChatCFD’s innovations include: (1) an interactive multi-modal input mechanism integrating PDFs, meshes, and dialogue with physical constraints, providing a ChatGPT-like interface for guided literature analysis, case selection, and parameter specification, reducing operational barriers; (2) specialized AI agents leveraging CFD expertise for task interpretation, feature library construction via OpenFOAM preprocessing, and iterative reflection-refinement using DeepSeek-R1 for reasoning and DeepSeek-V3 for precise modifications; and (3) a robust error reflection system with ReferenceRetriever (tutorial templates) and ContextRetriever (case configurations) for real-time error classification and resolution.
Experimental validations on 315 basic cases (205 benchmark tutorials, 110 perturbed variants) yield an 82.1% operational success rate, outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), with lower success in multiphase/combustion due to interdisciplinary complexity. On literature-derived cases, ChatCFD achieves 60–80% success, with ablation studies showing comprehensive interpretation boosts accurate configuration from 0% to 40% for NACA0012 and 30% for Nozzle, while ContextRetriever improves operational success from 0% to 60% in Nozzle by resolving coupled-physics errors. Turbulence model analyses reveal 40% success for common models (e.g., Spalart-Allmaras) vs. 10% for rare ones (e.g., RNG k-$\epsilon$), attributed to data scarcity. Physics coupling studies indicate compressible flows demand tripled resources due to interdependencies, and LLM bias toward incompressible setups causes dimensional errors in complex cases. These findings highlight ChatCFD’s robustness, efficiency, and cost-effectiveness, establishing benchmarks for AI-driven CFD.
ChatCFD effectively addresses the critical gaps in CFD automation agents. First, it integrates domain-specific knowledge through structured RAG and expert-designed agents, mitigating LLM training scarcity to enable precise boundary and initial condition specifications. Second, its modular, MCP-compatible design supports collaborative multi-agent networks, integrating specialized meshing agents to handle complex geometries beyond blockMesh. Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future. These advancements accelerate scientific discovery in fluid mechanics and engineering by automating hypothesis testing, parameter optimization, and multiphysics simulations.
ACKNOWLEDGEMENTS
This work is sponsored by the National Natural Science Foundation of China, Grant No. 92470127, and the Overseas Postdoctoral Talents Program in Guangdong Province.
AUTHOR DECLARATIONS
Conflict of Interest
The authors declare no conflicts of interest.
Author Contributions
E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft. Kang Hu: Methodology, Software, Data curation, Formal analysis, Resources. Zhuowen Wu: Methodology. Jiangyang Ge: Data curation, Formal analysis. Jiawei Miao: Software, Resources. Yuzhi Zhang: Conceptualization, Software, Resources. He Sun: Resources, Writing - Review. Weizong Wang: Resources, Writing - Review, Funding acquisition. Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
Data Availability
The data supporting the findings are available at https://github.com/ConMoo/ChatCFD.
Appendix A Appendix: detailed analysis of two literature-derived cases
Figure 10 presents a detailed analysis of the NACA0012 Case 1 performance across the five distinct Configurations (A through E), based on averaged results from 10-step operational success runs. This analysis yields several key insights into the system’s operational efficiency and cost-effectiveness.
The evolution of reflection iterations, illustrated in Figure 10(a), demonstrates an encouraging improvement pattern. As retrieve modules were integrated from configuration A to C, we observed a substantial decrease in required reflection iterations, followed by stabilization. This observation aligns with MetaOpenFOAM’s findings [36] regarding RAG positive impact on LLM response quality and CFD agent performance. The subsequent stability in reflection counts from configurations C to E suggests that additional RAG implementations, while beneficial for other aspects, had minimal impact on ChatCFD’s reflection behavior. This plateau is reasonable given that ContextRetriever and comprehensive paper interpretation modules were primarily designed to enhance performance on physical-coupled cases and setup accuracy respectively, rather than affecting reflection behaviors for weak coupling cases.
An analysis of token consumption and associated costs, depicted in Figure 10(b), reveals a pattern of improving resource utilization efficiency. Transitioning from Configuration A to D, while the token consumption of the DeepSeek-R1 model remained relatively stable, the DeepSeek-V3 model’s consumption decreased markedly. This reduction contributed to lower per-case operational costs (decreasing from \$0.078 to \$0.071). This cost improvement, concurrent with a significant increase in operational success rates from 20% to 80%, underscores the enhanced overall effectiveness of the ChatCFD system.
The transition from Configuration D to E resulted in a notable enhancement of system accuracy, as illustrated in Figure 6. This advancement was primarily achieved by improving the paper interpretation module’s efficacy in extracting CFD case setups during Stage 2 and by ensuring greater consistency of these setups throughout the file correction processes in Stage 3. Although this upgrade led to increased DeepSeek-R1 token consumption and a corresponding 30% rise in operational cost (from \$0.071 to \$0.091), the substantial improvement in accurate case configuration—from 0% to 40%—justifies this investment.
Figure 10(c) illustrates the token consumption patterns per reflection iteration for both DeepSeek-R1 and DeepSeek-V3 models across the different Configurations. From Configuration A to E, the DeepSeek-R1’s consumption increases to more than two times, where the increasing consumption are mostly due to the integration of the ContextRetriever module in Configuration C. The DeepSeek-V3’s consumption initially rose due to expanded error handling capabilities and reference case integration through ReferenceRetriever. The subsequent stabilization of DeepSeek-V3 token usage from Configuration C to E reflects the maturation of the reflection module’s design, achieving optimal operational efficiency without necessitating additional DeepSeek-V3 module invocations.
Figure 10: Results for NACA0012 Case 1 using Configurations A to E, showing averages of 10-step operational success runs. Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
Figure 11: Results for Nozzle Case 5 using Configurations D and E, showing averages of 10-step operational success runs. Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
Figure 11 presents a comparative performance analysis for the Nozzle Case 5, specifically contrasting Configurations D and E, based on simulations that successfully completed ten steps. Configurations A through C were omitted from this comparison due to their inability to consistently reach this ten-step benchmark. The key differentiator between Configurations D and E is the integration of the comprehensive paper interpretation module’s enhanced article interpretation capabilities in Configuration E. Notwithstanding this enhancement, Figure 11(a) indicates that the number of reflection iterations remained largely comparable between these two configurations, with only marginal reductions observed for Configuration E.
The deployment of comprehensive paper interpretation module resulted in increased token utilization, as depicted in Figure 11(b). Configuration E demonstrated higher token consumption by the DeepSeek-R1 model, whereas the DeepSeek-V3 model’s token usage remained stable across both Configurations D and E. In terms of operational cost, the average per-case expenditure saw a modest increase from \$0.28 to \$0.31, an approximate 10% increment. Nevertheless, this additional investment yielded significant improvements in performance, particularly in case setup accuracy, which, as shown in Figure 6, rose substantially from 0% to 30%.
Further examination, illustrated in Figure 11(c), reveals that the average token consumption per reflection iteration for both the DeepSeek-R1 and DeepSeek-V3 models was comparable across Configurations D and E. This consistency underscores the efficacy of the system’s architectural design, indicating that the comprehensive paper interpretation module detailed interpretation component operates as a distinct layer that augments overall system performance without adversely affecting the fundamental reflection mechanisms.
Table 2: Statistical averages of accurate configured cases for Naca0012 Case 1 and Nozzle Case 5.
Metrics Naca0012, Case 1 Nozzle, Case 5
Physical models Flow type: Incompressible Solver: simpleFoam Turbulence: Spalart-Allmaras Flow type: Compressible Solver: rhoCentralFoam Turbulence: Spalart-Allmaras Thermo model: hePsiThermo
Number of configuration files 9 12
Average tokens for all configuration files per case 1,792 2,647
Configuration files 0/p 0/U 0/nut 0/nuTilda constant/transportProperties constant/turbulenceProperties system/controlDict system/fvSchemes system/fvSolution 0/p 0/U 0/nut 0/nuTilda 0/T 0/alphat constant/transportProperties constant/turbulenceProperties constant/thermodynamicProperties system/controlDict system/fvSchemes system/fvSolution
DeepSeek-R1 call count 19.1 48.3
DeepSeek-V3 call count 17.4 35.2
Appendix B Appendix: CFD results of Cases 1 to 5
Figure 12: Velocity magnitude contours for NACA0012 Cases 1 to 4 at 10° angle of attack using four turbulence models. (a) Spalart-Allmaras model, (b) k-$\epsilon$ model, (c) k-$\omega$ SST model, (d) RNGk-$\epsilon$ model.
Figure 13: Comparison of pressure coefficients between CFD simulation and experimental data for NACA0012 airfoil at 10° angle of attack, using (a) Spalart-Allmaras model, (b) k-$\epsilon$ model, (c) k-$\omega$ SST model, and (d) RNGk-$\epsilon$ model. Symbols represent experimental results reported by Gregory et al. [37]
Figures 12 and 13 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models. As can be seen, the velocity contours from the four cases are similar, and the pressure coefficient distributions agree well with the experimental results [37].
Figure 14: (a) Experimental schlieren image [35] and (b) numerical schlieren image of Case 5 for the Nozzle case at nozzle pressure ratio = 3.
Figure 14 illustrates the experimental and numerical schlieren images for the nozzle case at a nozzle pressure ratio of 3. The experimental results are reported in Ref. [35]. The numerical results exhibit a strong correspondence with the experimental observations, accurately capturing key flow features. The figure clearly depicts the formation of a Mach stem, resulting from the intersection and reflection of shock waves from the ramp and flap. This interaction also reveals the characteristic Mach reflection, distinguished by its $\lambda$-shock wave structure.